# Reassessing quantum-thermodynamic enhancements in continuous thermal machines

### v2

@@ Appendix F Details on the NIC machine
For symmetric rates, we recover the situation where |$\ket{\beta}$ is completely decoupled from the states |$\ket{0}$ and |$\ket{1}$ ($\gamma_{\mathrm{h}}^{\beta}=0$), leading to a classical three-level system with local jumps induced by thermal baths at same temperatures than the original NIC machine, plus an extra uncoupled level (dark state). Considering Eqs. (38), (39) and (40), it becomes evident that in this limit, the quantum system and its classical analog become indistinguishable once the probability of the independent (dark) level is set to zero. On the other hand, if we approach this symmetric case asymptotically, the first inequality in Eq. (41) indicates that, in principle, the quantum system is more precise than the classical equivalent. However, it is important to note that in this limit, fluctuations in the system currents will diverge due to a first-order phase transition, as shown in Ref. [69]. That leads to $\mathcal{R}=0$ as can be appreciated in Fig. 10 for $\gamma_{\rm h}^{\rm a}=\gamma_{\rm h}^{\rm b}$, where ${\rm Var}[\dot{W}]$ diverges (vertical dotted line). The green and red areas in Fig. 10 represent respectively the regions of parameters where quantum advantage or disadvantages are obtained. This finding aligns with the observation that, in the symmetric case, the change of variables completely decouples coherences from populations. As mentioned before, in that case one can always redefine the relevant thermal machine under consideration by excluding the independent (dark) level. In that case, the remaining (connected) thermal machine will be exactly equal as their classical counterpart.

### v4

@@ Appendix H Details on the NIC machine analysis
For symmetric rates, we recover the situation where |$\ket{\beta}$ is completely decoupled from the states |$\ket{0}$ and |$\ket{1}$ ($\gamma_{\mathrm{h}}^{\beta}=0$), leading to a classical three-level system with local jumps induced by thermal baths at same temperatures than the original NIC machine, plus an extra uncoupled level (dark state). Considering Eqs. (22), (23) and (24), it becomes evident that in this limit, the quantum system and its classical analog become indistinguishable once the probability of the independent (dark) level is set to zero. On the other hand, if we approach this symmetric case asymptotically, the first inequality in Eq.  (25) indicates that, in principle, the quantum system is more precise than the classical equivalent. However, it is important to note that in this limit, fluctuations in the system currents will diverge due to a first-order phase transition, as shown in Ref. [60]. That leads to $\mathcal{R}=0$ as can be appreciated in Fig. 7 for $\gamma_{\rm h}^{\rm a}=\gamma_{\rm h}^{\rm b}$, where ${\rm Var}[\dot{W}]$ diverges (vertical dotted line). The green and red areas in Fig. 7 represent, respectively, the regions of parameters where quantum advantages or disadvantages are obtained. This finding aligns with the observation that, in the symmetric case, the change of variables completely decouples the coherences from the populations. As mentioned before, in that case one can always redefine the relevant thermal machine under consideration by excluding the independent (dark) level. In that case, the remaining (connected) thermal machine will be exactly equal as their classical counterpart.